# Efficient and Context-Aware Label Propagation for Zero-/Few-Shot Training-Free Adaptation of Vision-Language Model

### v1

@@ 3.5 Overall Algorithm
1:Input: Testing data stream $\mathcal{D}_{u}$, textual prototypes $\mathcal{P}$, few-shot features $\mathcal{D}_{l}$, and label propagation iterations $T$
@@
3:Initial graph $W=\mathbf{0}^{N_{p}+N_{l}\times N_{p}+N_{l}}$, initial label $Y^{0}_{p}$, $Y^{0}_{l}$ $Y_{u}^{0}$ by Eq. 2 # Initialization
@@
12:     $y_{i}=\arg\max_{c}Y_{ic}^{u}$# Testing sample label
@@ 4.1 Experimental Settings
Datasets: We selected a diverse range of datasets for evaluation, grouped into three main categories: fine-grained datasets, style-transfer datasets, and corruption datasets, covering a total of 30 test scenarios. Fine-Grained Datasets: We evaluated 11 widely recognized benchmarks, including ImageNet (Deng et al., 2009), Flowers102 (Nilsback & Zisserman, 2008), DTD (Cimpoi et al., 2014), OxfordPets (Parkhi et al., 2012), StanfordCars (Krause et al., 2013), UCF101 (Soomro, 2012), Caltech101 (Fei-Fei, 2004), Food101 (Fei-Fei, 2004), SUN397 (Xiao et al., 2010), FGVCAircraft (Maji et al., 2013), and EuroSAT (Helber et al., 2019). Style-Transfer Datasets: We used five benchmarks: ImageNet-V2 (Recht et al., 2019), ImageNet-A (Hendrycks et al., 2021b), ImageNet-R (Hendrycks et al., 2021a), and ImageNet-Sketch (Wang et al., 2019). Out-of-Distribution Datasets: We assessed 15 different domains from ImageNet-C (Hendrycks & Dietterich, 2019) with the highest severity 5, including Gaussian Noise, Shot Noise, Impulse Noise, Defocus Blur, Frosted Glass Blur, Motion Blur, Zoom Blur, Snow, Frost, Fog, Brightness, Contrast, Elastic Transform, Pixelate, and JPEG corruptions.
Implementation Details: For all experiments, we utilize the pretrained CLIP model (Radford et al., 2021) with both ResNet-based and ViT-based architectures, specifically ResNet50 (He et al., 2016) and ViT-B/16 (Dosovitskiy, 2020). Unlike previous works that tuned hyperparameters for each experiment, we simplify the process by using a consistent set of hyperparameters across all experiments. In constructing the KNN graph, each test image is connected to 3 text embeddings, 8 embeddings from the test memory bank, and 8 from few-shot images (if applicable). The exponential factor $\gamma$ for graph sparsification is set to 10, update factor $\beta$ for $\hat{Y}_{ui}$ to 0.2, factor $\alpha$ for label propagation to 1.0. For efficient inference, we limit the label propagation iterations $T$ to 3. We also adopt the dataset splits and textual prompts from Pratt et al. (2023); Zhang et al. (2022).
Competing Methods: We select a diverse set of competing methods for both zero-shot and few-shot adaptation scenarios to create a comprehensive benchmark. These methods can be broadly categorized into two groups: prompt-tuning methods and adapter-based methods. For prompt tuning, we include TPT (Shu et al., 2022), DiffTPT (Feng et al., 2023), CoOp (Zhou et al., 2022b), CoCoOp (Zhou et al., 2022a), among others. For adapter-based methods, we consider CLIP-Adapter Gao et al. (2024), TIP-Adapter Zhang et al. (2022), and the most recent approaches such as DMN (Zhang et al., 2024b), TDA (Karmanov et al., 2024), and ZLaP (Kalantidis et al., 2024), along with several other methods. Finally, we benchmark our proposed method, ECALP, on all tasks.
@@ 4.2 Evaluation of Zero-shot Adaptation
Fine-Grained Categorization Tasks: We present the results of zero-shot adaptation of CLIP model to fine-grained categorization downstream tasks in Tab. 1. We make the following observations from the results. i) Our proposed method consistently outperforms other methods across both architectures, showcasing robust generalization and strong performance, especially in the ViT-B/16 architecture. The ViT-based models generally perform better than their ResNet-based counterparts, reflecting the advantages of transformer-based architectures in handling diverse benchmarks. ii) DMN performs strongly across both architectures, particularly excelling in fine-grained classification tasks like Pets and Aircraft. However, the original DMN, denoted as DMN* in the table, implements exhaustive searching for the fusion hyper-parameters based on testing set for the individual downstream tasks and the performance drops by more than 1% when hyper-parameter searching is removed. Nevertheless, both DMN* and its variant, DMN with fixed hyper-parameters, are worse in general compared with us. Also, ZLaP requires additional unlabelled training set data to initialize the graph. iii) Some methods show strengths in specific benchmarks. For example, DiffPT excels in handling corruption datasets. The state-of-the-art label propagation based method, ZLap, achieves much lower performance compared with us. This is attributed to the more effective context-aware reweighting and iterative propagation strategies.
Table 1: Zero-shot adaptation of CLIP on fine-grained categorization downstream tasks. DMN* refers to the original method that exhaustively searches the fusion hyper-parameter with testing data ground-truth on each individual downstream task (not for direct comparison with other methods). ZLaP requires additional unlabelled training set data.
@@
Style-Transfer Tasks: We further evaluate the performance on style-transfer downstream tasks with results presented in Tab. 2. Our analysis reveals several key insights. i) Our proposed method consistently surpasses the baseline CLIP models, such as CLIP-RN50 and CLIP-ViT/B-16, with significant improvements across all datasets, especially in the averaged scores, where we observe a gain of approximately 24% for CLIP-RN50 and 7% for CLIP-ViT/B-16. This highlights the effectiveness of our approach in improving generalization without the need for additional training. ii) In comparison to state-of-the-art adaptation methods such as DMN and TDA, our method achieves higher accuracy averaged over all datasets, demonstrating the robustness of our method across different architectures. iii) Remarkably, even when compared to adaptation methods that require training, such as CoOp and CoCoOp, our method maintains a clear advantage, showcasing the superior performance of our method, even in scenarios where competing methods undergo task-specific fine-tuning.
@@
Method ImageNet-A ImageNet-V2 ImageNet-R ImageNet-S Mean CLIP-RN50 (Radford et al., 2021) 21.83 51.41 56.15 33.37 40.69 CoOp (Zhou et al., 2022b) 23.06 55.40 56.60 34.67 42.43 CoCoOp (Zhou et al., 2022a) 23.32 55.72 57.74 34.48 42.82 TPT (Shu et al., 2022) 26.67 54.70 59.11 35.09 43.89 DiffTPT (Feng et al., 2023) 31.06 55.80 58.80 37.10 45.69 CALIP (Guo et al., 2023) 23.96 53.70 60.81 35.61 43.52 TDA (Karmanov et al., 2024) 30.29 55.54 62.58 38.12 46.63 DMN (Zhang et al., 2024b) 28.57 56.12 61.44 39.84 46.49 ECALP (Ours) 28.71 56.91 63.69 41.50 47.70 CLIP-ViT-B/16 (Radford et al., 2021) 47.87 60.86 73.98 46.09 57.20 CoOp (Zhou et al., 2022b) 49.71 64.20 75.21 47.99 59.14 CoCoOp (Zhou et al., 2022a) 50.63 64.07 76.18 48.75 59.91 MaPLe (Khattak et al., 2023) 50.90 64.07 76.98 49.15 60.28 TPT (Shu et al., 2022) 54.77 63.45 77.06 47.94 60.81 DiffTPT (Feng et al., 2023) 55.68 65.10 75.00 46.80 60.65 TDA (Karmanov et al., 2024) 60.11 64.67 80.24 50.54 63.89 DMN (Zhang et al., 2024b) 58.28 65.17 78.55 53.20 63.80 ECALP (Ours) 58.33 65.76 80.73 54.70 64.88
Out-of-Distribution Tasks: We finally evaluate on adaptation to out-of-distribution downstream tasks. As seen from Tab. 3, our proposed method consistently outperforms state-of-the-art competing methods on almost all types of corruptions with both CNN and transformer backbones.
Table 3: Zero-shot adaptation of CLIP on corrupted downstream tasks. ZLaP requires additional unlabelled training set data.
@@ 4.3 Evaluation of Few-shot Adaptation
For few-shot adaptation tasks, we evaluate the adaptation performance on fine-grained categorization benchmarks, with [1, 2, 4, 8, 16] shots of each class, as presented in Fig.2. In line with (Silva-Rodriguez et al., 2024), we establish a fair benchmark by comparing existing methods using their default hyper-parameter values, as specified in the corresponding papers, without conducting a hyper-parameter search. We compare our method with several existing training-free approaches, including DMN(Zhang et al., 2024b), APE (Zhu et al., 2023), TIP-Adapter (Zhang et al., 2022), and TIP-X (Udandarao et al., 2023). Our method outperforms all competing methods on average across 11 datasets and consistently achieves superior results across most datasets and shot counts. Furthermore, it maintains high performance even in low-shot scenarios, underscoring its efficiency and robustness.
@@ 4.4 Ablation & Additional Study
Investigating Graph Construction: We further investigate the necessity of each subgraph of the overall graph. Specifically, we consider the three subgraphs including testing samples to prototypes $W_{up}$, within testing samples $W_{u}$ and testing samples to few-shot samples $W_{lu}$. As seen from Tab. 5, all three datasets benefit benefit substantially from having the subgraph connecting to textual prototypes, suggesting . When the subgraph within testing samples $W_{u}$ is included, the performance further improves, though more significantly on DTD and UCF, suggesting the manifold is helpful to the propagation of labels. Finally, including the testing sample subgraph $W_{u}$ is also helpful to the few-shot cases.
Computational Efficiency. To assess the efficiency of our proposed ECALP method, we measure the wall-clock time during the zero-shot adaptation task on fine-grained datasets, as shown in Tab. 6. All experiments are conducted using a single RTX 3090 GPU and an AMD EPYC 7302 CPU. ECALP, along with other training-free adaptation methods, does not significantly increase computational overhead and demonstrates at least a 30$\times$ speedup compared to approaches requiring training. Notably, ECALP achieves the best performance with low computational cost.
@@ 5 Conclusion
Adapting pre-trained vision language model for downstream tasks opens a new paradigm for crafting computer vision models. We aim to better exploit the observed unlabeled data, i.e. the data manifold and propose a unified method based on label propagation. In particular, we addressed the challenges of improving the computation efficiency of label propagation via iterative and incremental solution. We also proposed a simple dynamic graph expansion strategy to accommodate inductive inference. Furthermore, we notice that the similarity metric adopted by existing methods overlooked the diverse information captured by VLM and thus propose a context-aware edge reweighting based on the downstream task information. We carried out experiments on fine-grained categorization and out-distribution downstream tasks and achieved the state-of-the-art performance on all datasets.
@@ A.1.1 Static Graph Construction Complexity
$C_{S}=O\left(\sum_{n=1}^{N_{v}}n\log n\right)\approx O\left(\int_{1}^{N_{v}}x\log x,dx\right)=O\left(N_{v}^{2}\log N_{v}\right)$ (11)
@@ A.1.2 Dynamic Graph Expansion Complexity
A.1.3 Summary
In summary, constructing a static KNN graph has a total complexity of $O(N_{v}^{3}+N_{v}^{2}\log N_{v})$, which becomes computationally expensive as the size of the dataset increases. In contrast, our dynamic graph expansion method reduces the complexity to $O(d\cdot N_{v}^{2})$ by limiting distance computations to the current node and existing nodes in the graph, while using a sparse affinity matrix for efficient neighbor selection. This approach offers a significant reduction in computational cost, making it highly scalable and suitable for large datasets.
@@ A.2.1 KNN Connection Study
As illustrated in Fig. 3, our dynamic graph expansion enables ECALP to achieve a 6x speedup compared to the static graph construction method on the ImageNet dataset with CLIP-RN50.
@@ A.2.3 Generalization to different VLMs
The ability of ECALP to improve zero-shot adaptation extends beyond CLIP, showcasing its effectiveness across various Vision-Language Models (VLMs). As demonstrated in Table 1, applying ECALP to multiple VLMs results in a noticeable improvement in ImageNet accuracy. For instance, when applied to BLIP (Li et al., 2022), the base ImageNet accuracy of 53.40% is significantly enhanced to 58.16%. Similarly, in the case of BLIP v2 (Li et al., 2023), the accuracy increases from 41.22% to 43.72% with ECALP. A notable improvement is also observed with ALBEF (Li et al., 2021), where ECALP raises the accuracy from 36.15% to 40.31%. These results highlight the generalization ability of our approach, making it applicable across different state-of-the-art VLMs, consistently improving their performance.
@@ A.4 Failure Case Analysis for Few-Shot Adaptation
In this section, we delve into the reasons behind the suboptimal performance of our ECALP method on the OxfordPets and Food101 datasets under few-shot adaptation. We bring in other two datasets, Flowers102 and EuroSAT, for comparative studies. The testing sample features are projected into 2D space via t-SNE (van der Maaten & Hinton, 2008), as illustrated in Fig. 9. We make the following observations. The testing samples are more cluttered and mixed up in OxfordPets and Food101 than in Flowers102 and EuroSAT. There are more separated sub-classes, i.e. isolated clusters within each semantic class, for OxfordPets and Food101. This suggests having a few-shot labeled samples is less effective for OxfordPets and Food101 datasets. In certain cases, an inappropriate selection of few-shot samples may even bias the adaptation.

### v2

@@ 3.5 Overall Algorithm
1:Input: Test data stream $\mathcal{D}_{u}$, textual prototypes $\mathcal{P}$, few-shot features $\mathcal{D}_{l}$, and label propagation iterations $T$
@@
3:Initial graph $W=\mathbf{0}$, initial label $Y^{0}_{p}$, $Y^{0}_{l}$ $Y_{u}^{0}$ by Eq. 2 # Initialization
@@
12:     $y_{i}=\arg\max_{c}Y_{ic}^{u}$# Test sample label
@@ 4.1 Experimental Settings
Datasets: We selected a diverse range of datasets for thorough evaluation, grouped into three main categories: fine-grained datasets, style-transfer datasets, and out-of-distribution datasets, which cover a total of 30 specific test scenarios. Fine-Grained Datasets: We evaluated 11 widely recognized benchmarks, including ImageNet (Deng et al., 2009), Flowers102 (Nilsback & Zisserman, 2008), DTD (Cimpoi et al., 2014), OxfordPets (Parkhi et al., 2012), StanfordCars (Krause et al., 2013), UCF101 (Soomro, 2012), Caltech101 (Fei-Fei et al., 2004), Food101 (Fei-Fei et al., 2004), SUN397 (Xiao et al., 2010), FGVCAircraft (Maji et al., 2013), and EuroSAT (Helber et al., 2019). Style-Transfer Datasets: We used five benchmarks: ImageNet-V2 (Recht et al., 2019), ImageNet-A (Hendrycks et al., 2021b), ImageNet-R (Hendrycks et al., 2021a), and ImageNet-Sketch (Wang et al., 2019). Out-of-Distribution Datasets: We extensively assessed 15 unique different domains from ImageNet-C (Hendrycks & Dietterich, 2019) with the highest severity 5, including Gaussian Noise, Shot Noise, Impulse Noise, Defocus Blur, Frosted Glass Blur, Motion Blur, Zoom Blur, Snow, Frost, Fog, Brightness, Contrast, Elastic Transform, Pixelate, and JPEG corruptions.
Implementation Details: For all experiments, we use the pretrained CLIP model (Radford et al., 2021) with both ResNet-based and ViT-based architectures, specifically ResNet50 (He et al., 2016) and ViT-B/16 (Dosovitskiy, 2020). Unlike previous works, which tuned hyperparameters separately for each experiment, we simplify the process by using a consistent set of hyperparameters across all experiments. In constructing the KNN graph, each test image is connected to 3 text embeddings, 8 embeddings from the test memory bank, and 8 from few-shot images (if applicable). The exponential factor $\gamma$ for graph sparsification is set to 10, the update factor $\beta$ for $\hat{Y}_{ui}$ is set to 0.2, and the factor $\alpha$ for label propagation is set to 1.0. For efficient inference, we limit the label propagation iterations $T$ to 3. We also adopt the dataset splits and textual prompts from Pratt et al. (2023); Zhang et al. (2022).
Competing Methods: We selected a diverse set of competing methods for both zero-shot and few-shot adaptation scenarios to create a comprehensive benchmark. These methods can be broadly categorized into two groups: prompt-tuning methods and adapter-based methods. For prompt tuning, we include TPT (Shu et al., 2022), DiffTPT (Feng et al., 2023), CoOp (Zhou et al., 2022b), CoCoOp (Zhou et al., 2022a), among others. For adapter-based methods, we consider CLIP-Adapter (Gao et al., 2024), TIP-Adapter (Zhang et al., 2022), and the most recent approaches such as DMN (Zhang et al., 2024b), TDA (Karmanov et al., 2024), and ZLaP (Kalantidis et al., 2024), along with several other methods. Finally, we benchmark our proposed method, ECALP, on all tasks.
@@ 4.2 Evaluation of Zero-shot Adaptation
Fine-Grained Categorization Tasks: We present the results of zero-shot adaptation of CLIP model to fine-grained categorization downstream tasks in Tab. 1. We make the following observations from the results. i) Our proposed method consistently outperforms other methods across both architectures, showcasing robust generalization and strong performance, especially in the ViT-B/16 architecture. The ViT-based models generally perform better than their ResNet-based counterparts, reflecting the advantages of transformer-based architectures in handling diverse benchmarks. ii) DMN performs strongly across both architectures, particularly excelling in fine-grained classification tasks like Pets and Aircraft. However, the original DMN, denoted as DMN* in the table, implements exhaustive searching for the fusion hyperparameters based on testing set for the individual downstream tasks and the performance drops by more than 1% when hyperparameter searching is removed. Nevertheless, both DMN* and its variant, DMN with fixed hyperparameters, are worse in general compared with us. Also, ZLaP requires additional unlabelled training set data to initialize the graph. iii) Some methods show strengths in specific benchmarks. For example, DiffPT excels in handling corruption datasets. The state-of-the-art label propagation based method, ZLap, achieves much lower performance compared with us. This is attributed to the more effective context-aware reweighting and iterative propagation strategies.
Table 1: Zero-shot adaptation of CLIP on fine-grained categorization downstream tasks. DMN* refers to the original method that exhaustively searches the fusion hyperparameter with testing data ground-truth on each individual downstream task (not for direct comparison with other methods). ZLaP requires additional unlabelled training set data.
@@
Style-Transfer Tasks: We further evaluate the performance on style-transfer downstream tasks with results presented in Tab. 2. Our analysis reveals several key insights. i) Our proposed method consistently surpasses the baseline CLIP models, such as CLIP-RN50 and CLIP-ViT/B-16, with significant improvements across all datasets, especially in the averaged scores, with an average gain of approximately 7%. This highlights the effectiveness of our approach in improving generalization without the need for additional training. ii) In comparison to state-of-the-art adaptation methods such as DMN and TDA, our method achieves higher accuracy averaged over all datasets, demonstrating that our method is robust across different architectures. iii) Remarkably, even when compared to adaptation methods that require training, such as CoOp and CoCoOp, our method maintains a clear advantage, showcasing the superior performance of our method, even in scenarios where competing methods undergo task-specific fine-tuning.
@@
Method ImageNet-A ImageNet-V2 ImageNet-R ImageNet-S Mean CLIP-RN50 (Radford et al., 2021) 21.83 51.41 56.15 33.37 40.69 CoOp (Zhou et al., 2022b) 23.06 55.40 56.60 34.67 42.43 CoCoOp (Zhou et al., 2022a) 23.32 55.72 57.74 34.48 42.82 TPT (Shu et al., 2022) 26.67 54.70 59.11 35.09 43.89 DiffTPT (Feng et al., 2023) 31.06 55.80 58.80 37.10 45.69 CALIP (Guo et al., 2023) 23.96 53.70 60.81 35.61 43.52 TDA (Karmanov et al., 2024) 30.29 55.54 62.58 38.12 46.63 DMN (Zhang et al., 2024b) 28.57 56.12 61.44 39.84 46.49 ECALP (Ours) 28.80 56.92 63.68 41.51 47.73 CLIP-ViT-B/16 (Radford et al., 2021) 47.87 60.86 73.98 46.09 57.20 CoOp (Zhou et al., 2022b) 49.71 64.20 75.21 47.99 59.14 CoCoOp (Zhou et al., 2022a) 50.63 64.07 76.18 48.75 59.91 MaPLe (Khattak et al., 2023) 50.90 64.07 76.98 49.15 60.28 TPT (Shu et al., 2022) 54.77 63.45 77.06 47.94 60.81 DiffTPT (Feng et al., 2023) 55.68 65.10 75.00 46.80 60.65 TDA (Karmanov et al., 2024) 60.11 64.67 80.24 50.54 63.89 DMN (Zhang et al., 2024b) 58.28 65.17 78.55 53.20 63.80 ECALP (Ours) 58.52 65.72 80.77 54.66 64.92
Out-of-Distribution Tasks: Finally, we evaluate the model’s adaptation to out-of-distribution downstream tasks. As seen from Tab. 3, our proposed method consistently outperforms state-of-the-art competing methods on almost all types of corruptions with both CNN and transformer backbones.
Table 3: Zero-shot adaptation of CLIP on out-of-distribution downstream tasks. ZLaP requires additional unlabelled training set data.
@@ 4.3 Evaluation of Few-shot Adaptation
For few-shot adaptation tasks, we evaluate the adaptation performance on fine-grained categorization benchmarks, with [1, 2, 4, 8, 16] shots of each class, as presented in Fig. 2. In line with (Silva-Rodriguez et al., 2024), we establish a fair benchmark by comparing existing methods using their default hyperparameter values, as specified in the corresponding papers, without conducting a hyperparameter search. We compare our method with several existing training-free approaches, including DMN(Zhang et al., 2024b), APE (Zhu et al., 2023), TIP-Adapter (Zhang et al., 2022), and TIP-X (Udandarao et al., 2023). Our method outperforms all competing methods on average across 11 datasets and consistently achieves superior results across most datasets and shot counts. Furthermore, it maintains high performance even in low-shot scenarios, underscoring its efficiency and robustness.
@@ 4.4 Ablation & Additional Study
Investigating Graph Construction: We further investigate the necessity of each subgraph of the overall graph. Specifically, we consider the three subgraphs including test samples to prototypes $W_{up}$, within test samples $W_{u}$ and test samples to few-shot samples $W_{lu}$. As seen from Tab. 5, all three datasets benefit substantially from having the subgraph connecting to textual prototypes. When the subgraph within test samples $W_{u}$ is included, the performance further improves, though more significantly on DTD and UCF, suggesting the manifold is helpful to the propagation of labels. Finally, including the test sample subgraph $W_{u}$ is also helpful to the few-shot cases.
Computational Efficiency. To assess the computational efficiency of our proposed ECALP method, we measure the wall-clock time during the zero-shot adaptation task on fine-grained datasets, as shown in Tab. 6. All experiments are conducted using a single RTX 3090 GPU and an AMD EPYC 7302 CPU. ECALP, along with other training-free adaptation methods, does not significantly increase computational overhead and demonstrates at least a 30$\times$ speedup compared to approaches requiring training. Notably, ECALP achieves the best performance with low computational cost.
@@ 5 Conclusion
Adapting pre-trained vision language model for downstream tasks introduces a new paradigm for crafting computer vision models. We aim to better exploit the observed unlabeled data, i.e. the data manifold and propose a unified method based on label propagation. In particular, we addressed the challenges of improving the computation efficiency of label propagation via iterative and incremental solution. We also proposed a simple dynamic graph expansion strategy to accommodate inductive inference. Furthermore, we observe that the similarity metric adopted by existing methods overlooked the diverse information captured by VLM and thus propose a context-aware edge reweighting based on the downstream task information. We carried out experiments on fine-grained categorization and out-distribution downstream tasks and achieved the state-of-the-art performance on all datasets.
Acknowledgements
This work was supported by the National Natural Science Foundation of China (NSFC) under Grant 62106078; by the Agency for Science, Technology and Research, Singapore (A*STAR) under Grant M23L7b0021; by the Guangdong Provincial Key Laboratory of Human Digital Twin under Grant 2022B1212010004; by the National Natural Science Foundation of China (NSFC) under Grant 61701181; by the Guangdong R&D key project of China under Grant 2019B010155001.
@@ A.1.1 Static Graph Construction Complexity
$C_{S}=O\left(\sum_{n=1}^{N_{v}}n\log n\right)\approx O\left(\int_{1}^{N_{v}}x\log x\,dx\right)=O\left(N_{v}^{2}\log N_{v}\right)$ (11)
@@ A.2.1 KNN Connection Study
As illustrated in Fig. 3, our dynamic graph expansion enables ECALP to achieve an approximate 6$\times$ speedup compared to the static graph construction method on the ImageNet dataset with CLIP-RN50.
@@ A.2.3 Generalization to different VLMs
The ability of ECALP to improve zero-shot adaptation extends beyond CLIP, showcasing its effectiveness across various Vision-Language Models (VLMs). As demonstrated in Tab. 8, applying ECALP to multiple VLMs results in a noticeable improvement in ImageNet accuracy. For instance, when applied to BLIP (Li et al., 2022), the base ImageNet accuracy of 53.40% is significantly enhanced to 58.16%. Similarly, in the case of BLIP v2 (Li et al., 2023), the accuracy increases from 41.22% to 43.72% with ECALP. A notable improvement is also observed with ALBEF (Li et al., 2021), where ECALP raises the accuracy from 36.15% to 40.31%. These results highlight the generalization ability of our approach, making it applicable across different state-of-the-art VLMs, consistently improving their performance.
@@ A.4 Failure Case Analysis for Few-Shot Adaptation
In this section, we delve into the reasons behind the suboptimal performance of our ECALP method on the OxfordPets and Food101 datasets under few-shot adaptation. We bring in other two datasets, Flowers102 and EuroSAT, for comparative studies. The test sample features are projected into 2D space via t-SNE (van der Maaten & Hinton, 2008), as illustrated in Fig. 9. We make the following observations. The test samples are more cluttered and mixed up in OxfordPets and Food101 than in Flowers102 and EuroSAT. There are more separated sub-classes, i.e. isolated clusters within each semantic class, for OxfordPets and Food101. This suggests having a few-shot labeled samples is less effective for OxfordPets and Food101 datasets. In certain cases, an inappropriate selection of few-shot samples may even bias the adaptation.